Problem 2 The fuel-cost function in $/h of two thermal plants are ?1 = 320 + 6.2?1 + 0.004?1 2 , ?2 = 200 + 6.0?2 + 0.003?2 2 , where ?1 and ?2 are in MW. Plant outputs are subject to the following limits (in MW)
Answer #1

The fuel cost function m &th of huo humal plank au. 9=320+612 P140-004 PM given as - (z = 200 +6.0l2t0.003022. power loss. PeTo calculate gali) 1. We need. 19 0.00420.000125x6.2 0.003+0.0000 625 X6 4P70X0.00 0125) 260.00 34 7 X0-0000625)** = 243. 27.